OpenStudy (anonymous):

how do you simplify √-4 using the imaginary number i ?

OpenStudy (turingtest):

\[\sqrt{-4}=\sqrt{-1\cdot4}=\sqrt{-1}\cdot\sqrt4\]you should be able to do both those square roots individually. multiply the answers

OpenStudy (turingtest):

\[\sqrt{-1}=i\]